### v3

# arXiv:2512.17131v3[cs.LG] 27 Feb 2026

Abstract
We propose Generalized Primal Averaging (GPA), an extension of Nesterov’s method that unifies and generalizes recent averaging-based optimizers like single-worker DiLoCo and Schedule-Free, within a non-distributed setting. While DiLoCo relies on a memory-intensive two-loop structure to periodically aggregate pseudo-gradients using Nesterov momentum, GPA eliminates this complexity by decoupling Nesterov’s interpolation constants to enable smooth iterate averaging at every step. Structurally, GPA resembles Schedule-Free but replaces uniform averaging with exponential moving averaging. Empirically, GPA consistently outperforms single-worker DiLoCo and AdamW with reduced memory overhead. GPA achieves speedups of 8.71%, 10.13%, and 9.58% over the AdamW baseline in terms of steps to reach target validation loss for Llama-160M, 1B, and 8B models, respectively. Similarly, on the ImageNet ViT workload, GPA achieves speedups of 7% and 25.5% in the small and large batch settings respectively. Furthermore, we prove that for any base optimizer with $\mathcal{O}(\sqrt{T})$ regret, where $T$ is the number of iterations, GPA matches or exceeds the original convergence guarantees depending on the interpolation constants.
[…]
1 Introduction
As large language models (LLMs) demonstrate increasingly remarkable capabilities at scale (achiam2023gpt; grattafiori2024llama3herdmodels; liu2024deepseek), the pre-training phase has become one of the most resource intensive stages in the language model training pipeline. This has encouraged the development of training algorithms and optimizers that enhance the efficiency, scalability, and robustness of language model pre-training. One significant area of research is the design of training algorithms for scalable distributed learning. In this area, the DiLoCo algorithm has emerged as the leading practical approach (diloco; liu2024asynchronous; douillard2025streaming; charles2025communication).
Despite its practical success, the underlying reasons for DiLoCo’s effectiveness remain poorly understood. Importantly, DiLoCo is not limited to distributed training: single-worker DiLoCo outperforms AdamW even in the non-distributed setting. kallusky2025snoo suggest that this is due to its novel combination of the Nesterov optimizer with the Lookahead method (lookahead), called Step-$K$ Nesterov. The method accumulates multiple updates from a base optimizer on an inner set of weights, forming what is called a pseudo-gradient. It then applies Nesterov momentum to the pseudo-gradient to update an outer set of weights, and subsequently resets the inner weights to match the new values of the outer weights. On a 160 million parameter Llama model, single-worker DiLoCo achieves speedups of up to 6.32% in terms of steps to reach the final validation loss by AdamW; see Figure 1(b).
(a) Both GPA and single-worker DiLoCo, when using AdamW as their base optimizer, outperform the tuned AdamW baseline for training a 160M parameter Llama model. Notably, increasing the number of inner steps (up to 64) improves the performance of single-worker DiLoCo. GPA instead updates the parameters at every step using a heuristic to choose interpolation constants that approximately match the number of inner steps for single-worker DiLoCo.
(b) Speedup achieved by single-worker DiLoCo and GPA measured in terms of reduction in number of steps required to attain the final validation loss obtained by AdamW, across different effective numbers of inner steps. GPA attains the highest speedup of 8.71% when the effective inner steps is equal to 32. Single-worker DiLoCo does not exhibit any speedup for 128 inner steps.
Figure 1: Comparison of validation loss and speedup for AdamW, single-worker DiLoCo, and GPA on Llama-160M model. See Appendix 11.2 for similar results on Llama-1B.
Intriguingly, DiLoCo’s performance initially improves as the number of inner steps increases. With each inner step, DiLoCo’s outer weights drift farther from its inner weights, similar to meta-learning optimizers such as Reptile (nichol2018reptile) and First-Order MAML (finn2017model). As a result, updates to the outer weights occur only at periodic intervals, causing information from the data to be integrated in a discontinuous, choppy manner rather than smoothly at every iteration. This restriction on information flow to the outer weights appears unnecessary from an optimization perspective, yet counterintuitively improves its performance; see Figure 1(a).
Concurrently, the Schedule-Free optimizer recently won the AlgoPerf Algorithmic Efficiency challenge self-tuning track (Dahl2023AlgoPerf; schedule-free). Its core novelty lies in computing gradients at a point that interpolates between the uniform average of past weights and the current weights. Empirically, Schedule-Free matches the performance obtained by using learning rate schedules without using any schedule explicitly, while providing stronger theoretical last-iterate convergence guarantees similar to Polyak-Ruppert averaging (ruppert; polyak; polyak1992acceleration). However, its reliance on uniform averaging limits its flexibility and performance in some settings.
In this paper, we argue that these two lines of work – DiLoCo and Schedule-Free – are closely related and can be generalized and improved through a unified framework of primal averaging. Specifically, our contributions are as follows:
  • We propose a novel generalization of Nesterov’s method in its primal averaging formulation called Generalized Primal Averaging (GPA). The method can be interpreted as a smoothed version of single-worker DiLoCo that incrementally averages iterates at every step. It can also be viewed as a subtle change of Schedule-Free that replaces uniform averaging with exponential moving averaging through a decoupled interpolation parameter to improve its practical performance.
  • In contrast to single-worker DiLoCo, GPA eliminates the two-loop structure, thereby requiring only a single additional buffer with one less hyperparameter to tune. Because it incrementally averages iterates at every step, the method consistently exhibits more stable training behavior than single-worker DiLoCo.
  • Our experiments demonstrate that GPA consistently outperforms single-worker DiLoCo and AdamW on dense 160 million and 1 billion parameter language models. We validate our results on different modalities through an 8 billion parameter Llama code generation model and through a vision workload using ImageNet ViT in both small- and large-batch settings. In particular, on the Llama-160M, 1B, and 8B models, we find that GPA provides speedups of 8.71%, 10.13%, and 9.58%, respectively, in terms of steps to reach the baseline validation loss. Likewise, GPA obtains speedups of 7% and 25.5% (small and large batch) on the ImageNet ViT workloads.
  • We provide theoretical justification for GPA by proving convergence guarantees that show improved convergence over the base optimizer on stochastic convex and non-smooth functions under some settings.
2 Background
We frame language model pre-training as the expected risk minimization problem
$\min_{x\in\mathbb{R}^{n}}F(x)=\mathbb{E}_{\xi\sim\mathcal{D}}\left[f(x;\xi)\right],$ (1)
where $\xi\sim\mathcal{D}$ is drawn from an underlying stationary data distribution $\mathcal{D}$. We assume that each optimizer step has access to the stochastic minibatch gradient $g(x^{(t)};\xi^{(t)})\in\partial f(x^{(t)};\xi^{(t)})$ evaluated at each iteration $t$ on a minibatch of data $\xi^{(t)}$, over a total of $T$ steps. [^1] We assume that $f$ is convex for the convergence analysis, but we verify its performance on non-convex, possibly non-smooth functions.
We also assume that the base optimizer is of the form $x^{(t+1)}=x^{(t)}+\gamma^{(t)}d^{(t)}$ with learning rate $\gamma^{(t)}>0$ and search direction $d^{(t)}\in\mathbb{R}^{n}$. The search direction is most commonly defined as $d^{(t)}=-H^{(t)}m^{(t)}$, where $m^{(t)}\in\mathbb{R}^{n}$ is a gradient estimator, and $H^{(t)}\in\mathbb{R}^{n\times n}$ is a symmetric positive definite preconditioner matrix. This includes popular methods such as SGD, Adam, Shampoo, SOAP, AdEMAMix, or Muon for different choices of $m^{(t)}$ and $H^{(t)}$ (robbins1951stochastic; kingma2014adam; pmlr-v80-gupta18a; loshchilov2018decoupled; anil2020scalable; shi2023distributed; vyas2024soap; jordan2024muon; pagliardini2025the; eschenhagen2025purifying).
2.1 Different Formulations of Nesterov Momentum
Nesterov momentum has played a critical role in optimization for deep learning (pmlr-v28-sutskever13). Despite its importance, there is still substantial confusion in the literature regarding Nesterov’s formulation, as it can be written in at least seven different ways (adefazio-curvedgeom2019). These formulations are equivalent in the sense that a direct mapping exists between them, but they may not return the same iterate.
For instance, Nesterov’s method was popularized for deep learning in Sutskever’s formulation (pmlr-v28-sutskever13), which presents the algorithm as:
$\begin{split}b^{(t)}&=\mu b^{(t-1)}-\gamma^{(t)}g(x^{(t)}+\mu b^{(t-1)};\xi^{(t)}),\\ x^{(t+1)}&=x^{(t)}+b^{(t)},\end{split}$ (2)
where $\mu>0$ is the momentum hyperparameter and $b^{(t)}\in\mathbb{R}^{n}$ is the momentum buffer initialized at $b^{(0)}=0$. An alternative formulation, which we call the modern formulation, is used by software libraries such as PyTorch [^2] PyTorch SGD documentation and JAX [^3] Optax (JAX) SGD documentation due to its ease of use:
$\begin{split}b^{(t)}&=\mu b^{(t-1)}+g(x^{(t)};\xi^{(t)}),\\ x^{(t+1)}&=x^{(t)}-\gamma^{(t)}[\mu b^{(t)}+g(x^{(t)};\xi^{(t)})].\end{split}$ (3)
In both formulations, we maintain a momentum buffer that averages the gradients seen throughout the training process. However, unlike Sutskever’s formulation (Equation˜2), the modern formulation (Equation˜3) uses the iterate $x^{(t)}$ directly for the gradient computation, rather than the ancillary point $x^{(t)}+\mu b^{(t-1)}$, simplifying its practical implementation. If both formulations are run side-by-side with the same seed, they will evaluate gradients at exactly the same points, but their validation losses at iterates $x^{(t)}$ for each method will differ.
Our approach instead builds upon a third form, which we call the primal averaging formulation:
$\begin{split}y^{(t)}&=\mu x^{(t)}+(1-\mu)z^{(t)},\\ z^{(t+1)}&=z^{(t)}-\gamma^{(t)}g(y^{(t)};\xi^{(t)}),\\ x^{(t+1)}&=\mu x^{(t)}+\left(1-\mu\right)z^{(t+1)},\end{split}$ (4)
with $\mu\in[0,1)$. The first mention of this three-sequence form that we are aware of is by lanaccel, although it was only studied under a time-varying $\mu$.
Unlike the Sutskever and modern formulations framed in equations 2 and 3, the primal averaging formulation in Equation˜4 explicitly names two iterate sequences: a sequence where the gradients (or, more generally, the search directions) are computed at, i.e., the gradient computation sequence $\{y^{(t)}\}_{t=1}^{T}$, as well as another sequence used for model evaluation that accumulates a running average of updated iterates $\{z^{(t)}\}_{t=1}^{T}$, i.e., the model evaluation sequence $\{x^{(t)}\}_{t=1}^{T}$. Since $y^{(t)}$ interpolates the smoothed sequence $x^{(t)}$ and unsmoothed sequence $z^{(t)}$, it increases the contribution of the gradient update to $y^{(t)}$ compared to $x^{(t)}$. This explicit formulation is convenient for implementation and theoretical analysis, and naturally leads to a view of acceleration as built upon iterate averaging, rather than from the physics-inspired intuition of gradient averaging behind momentum that is more commonly introduced.
We summarize the relationship between the modern and primal averaging formulations in Proposition 1 below.
Proposition 1.
Given fixed learning rates $\gamma_{\operatorname{primal}},\gamma_{\operatorname{modern}}>0$, the primal averaging formulation of Nesterov’s method (Equation˜4) is equivalent to its modern formulation (Equation˜3) in the sense that
$y_{\operatorname{primal}}^{(t)}$ $=x_{\operatorname{modern}}^{(t)}\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{and}$ (5)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ b_{\operatorname{modern}}^{(t)}$ $=\frac{1}{\left(1-\mu\right)\gamma_{\operatorname{primal}}}\left(x_{\operatorname{primal}}^{(t)}-x_{\operatorname{primal}}^{(t+1)}\right),$
when $\mu_{\operatorname{primal}}=\mu_{\operatorname{modern}}=\mu$ and $\left(1-\mu\right)\gamma_{\operatorname{primal}}=\gamma_{\operatorname{modern}}$.
The proof of this simple statement is rather technical, so we defer it to Appendix 10. Similar formulations and equivalences can be derived for Polyak momentum (polyak1964some; defazio2020mom; ziyin2020laprop); see Appendix 8. We summarize these formulations and their differences in Table˜1.
Remark. It is important to acknowledge that the equivalence between the primal averaging and modern formulations of Nesterov momentum holds only when the learning rates are constant. When learning rate schedules are introduced, achieving this equivalence would require the momentum parameter to vary with each iteration. Furthermore, the restriction on the choice of $\mu$ differs between the modern and primal averaging formulations. These different interpretations based on gradient averaging versus iterate averaging produce different perspectives for hyperparameter tuning, which can have a significant impact on the algorithm’s practical performance.
Table 1: Summary of Nesterov, Schedule-Free, and GPA formulations. Here, $\mu$ is the momentum parameter, $\gamma^{(t)}$ is the learning rate, and $g(\cdot;\xi^{(t)})$ denotes the stochastic gradient. For a summary of Polyak momentum formulations, please refer to Appendix 8.
Formulation Update Equations Notes
[…]
2.2 Single-Worker DiLoCo and its Weaknesses
DiLoCo was originally introduced as a distributed algorithm for cross-datacenter training (diloco). The method computes multiple inner steps of a base optimizer on the inner weights, then applies Nesterov (Equation˜3) on the average pseudo-gradient, the difference between the previous and updated inner model weights, to update the outer weights. The inner weights are then reset to outer weights.
DiLoCo requires storing two additional optimizer states of the same shape as the model parameters: the momentum buffer $b^{(t)}$ and the current model parameters $x^{(t)}$ (also known as the outer weights). DiLoCo’s handling of fast inner weights and slow outer weights can be interpreted as a modified Lookahead method that applies Nesterov momentum to the outer weight updates (lookahead). The method was recently analyzed in khaled2025understanding, and demonstrated significant compute factor gains in the non-distributed setting in kallusky2025snoo. A simplified version of non-distributed or single-worker DiLoCo with $H$ inner steps of the base optimizer can be described as:
$p^{(t)}$ $=x^{(t)}-\operatorname{\texttt{BaseOptIteration}}(x^{(t)};\{\gamma^{(j)}\}_{j=1}^{H},H)$
$b^{(t)}$ $=\mu b^{(t-1)}+p^{(t)}$ (6)
$x^{(t+1)}$ $=x^{(t)}-\tilde{\gamma}[\mu b^{(t)}+p^{(t)}],$
where $\tilde{\gamma}>0$ is the outer learning rate and $\operatorname{\texttt{BaseOptIteration}}$ applies $H$ inner steps of the base optimizer to the iterate $x^{(t)}$ with inner learning rates $\{\gamma^{(j)}\}_{j=1}^{H}$. While DiLoCo originally introduced AdamW as the base optimizer, DiLoCo has been generalized to other optimizers such as Muon (therien2025muloco). A complete description of the algorithm is provided in Appendix 9. As noted in kallusky2025snoo, applying Nesterov on the pseudo-gradient with multiple inner steps is capable of surpassing the performance of the base optimizer alone, which explains DiLoCo’s ability to match the synchronous baseline, such as AdamW, in the multi-worker setting.
Weaknesses in DiLoCo’s hierarchical framework.
However, this two-level structure is undesirable. From an algorithmic perspective, one would prefer to average iterates on-the-fly, as opposed to averaging trajectories that implicitly contain multiple iterations of the base optimizer. From the users’ perspective, the two-level structure introduces an additional copy of the model weights required to compute the pseudo-gradient, and introduces additional hyperparameters to tune, e.g., the inner and outer learning rates, momentum, and number of inner steps. Lastly, from the distributed training perspective, DiLoCo couples the number of inner steps as a hyperparameter for both local SGD as well as for the modified Nesterov algorithm, causing the algorithm’s performance to counterintuitively improve as the number of inner steps increases. One would instead expect that communicating more often should always be beneficial. These challenges motivate the development of a new algorithm that removes the two-level structure while offering a separate hyperparameter that can smoothly average the observed iterates at every iteration.
2.3 Schedule-Free Learning
In parallel, Schedule-Free learning (SF) (schedule-free) was recently proposed as a wrapper to any base optimizer using a variant of the primal averaging formulation of Nesterov’s method (Equation˜4) for hyperparameter-free learning:
$\begin{split}y^{(t)}&=\mu x^{(t)}+(1-\mu)z^{(t)}\\ z^{(t+1)}&=z^{(t)}-\gamma g(y^{(t)};\xi^{(t)})\\ x^{(t+1)}&=\frac{t}{t+1}x^{(t)}+\Bigl(1-\frac{t}{t+1}\Bigr)z^{(t+1)}.\end{split}$ (7)
Originally designed to eliminate the need for manually specified learning rate schedules, Schedule-Free has demonstrated the surprising ability to not only match, but even surpass the practical performance of the original base optimizer. This is done by decoupling the momentum hyperparameter used in the $x^{(t)}$ and $y^{(t)}$ sequences, unlike the standard primal averaging formulation of Nesterov (Equation˜4). Through the choice of $\mu$, the method interpolates between uniform Polyak-Ruppert averaging and stochastic primal averaging (ruppert; polyak; tao2018).
Ignoring the hyperparameter-free learning problem, one could alternatively replace uniform averaging with exponential moving averaging of the iterates, which is commonly used in practice (morales2024exponential). This alternative suggests a different generalization of Nesterov momentum that may offer the potential flexibility necessary to reproduce DiLoCo’s convergence gains without the two-level structure.
3 Generalized Primal Averaging (GPA)
By decoupling the constants for the model evaluation and gradient computation sequences in the primal averaging formulation of Nesterov’s method (Equation˜4) and leveraging the observation of using exponential moving averaging in lieu of uniform averaging in Schedule-Free (Equation˜7), we introduce the Generalized Primal Averaging (GPA) framework:
$\begin{split}y^{(t)}&=\mu_{y}x^{(t)}+(1-\mu_{y})z^{(t)}\\ z^{(t+1)}&=z^{(t)}-\gamma^{(t)}g(y^{(t)};\xi^{(t)})\\ x^{(t+1)}&=\mu_{x}x^{(t)}+\left(1-\mu_{x}\right)z^{(t+1)}.\end{split}$ (8)
Here, $\mu_{x}\in[0,1)$ and $\mu_{y}\in[0,1]$ are independent hyperparameters that separately control the degree of interpolation used to maintain the model evaluation sequence $x^{(t)}$ and gradient computation sequence $y^{(t)}$. The additional hyperparameter $\mu_{x}$ serves as a smoothing or exponential moving average parameter that replaces Polyak-Ruppert averaging in Schedule-Free, while $\mu_{y}$ controls the amount of information flow into $y^{(t)}$. The complete pseudocode for a general base optimizer is provided in Algorithm 1.
Unlike the modern formulation of Nesterov momentum (Equation˜3) or DiLoCo (Equation˜6) built on (pseudo-)gradient averaging, GPA is defined based on the primal or iterate averaging framework. We argue that this provides a more meaningful characterization of the method. For example, the primal averaging interpretation naturally extends to other search directions by replacing $g(y^{(t)};\xi^{(t)})$ with the search direction $d^{(t)}$ evaluated at $y^{(t)}$. This extension is not intuitive from the gradient averaging perspective, as it would translate to averaging search directions (with potentially different, evolving preconditioners) in the momentum buffer.
Learning rate schedules. By replacing Polyak-Ruppert averaging with exponential moving averaging, GPA is not inherently schedule-free and requires the use of a learning rate schedule. To see why, observe that Polyak averaging places increasingly less weight $1/(t+1)$ on the most recent iterate $z^{(t+1)}$, which plays a similar role to learning rate scheduling (sandler2023trainingtrajectories; schedule-free). GPA instead places a constant weight $\mu_{x}$ on the most recent iterate $z^{(t+1)}$ by leveraging an exponential moving average, thereby requiring a learning rate schedule compared to Schedule-Free. This is reflected theoretically in their average-iterate versus last-iterate convergence properties.
Degenerate cases. The choice of $\mu_{x}$ and $\mu_{y}$ enables GPA to recover different averaging methods:
  • When $\mu_{y}=1$, $x^{(t)}=y^{(t)}$ and we recover stochastic primal averaging, or equivalently, LaProp (defazio2020mom; ziyin2020laprop); see Appendix 9.
  • When $\mu_{y}=0$, $x^{(t)}$ and $z^{(t)}=y^{(t)}$ become decoupled and we recover exponential moving averaging of the iterates (morales2024exponential).
  • When $\mu_{x}=0$, $x^{(t)}=y^{(t)}=z^{(t)}$ for any choice of $\mu_{y}$, and GPA reverts to the base optimizer.
Memory-efficient implementation. When $\mu_{y}>0$, GPA can be implemented with only one extra copy of the model weights – specifically, by storing $y^{(t)}$ and reconstructing $x^{(t)}$ from $y^{(t)}$ and $z^{(t)}$ during evaluation – unlike DiLoCo, which demands more memory overhead due to the global parameter and momentum buffers. The pseudocode is provided in Appendix 9.
Other properties. Because $\mu_{x},\mu_{y}\in[0,1]$, GPA preserves modular norm bounds of the model parameters (bernstein2024modular). More details on this property is provided in Appendix 9. GPA can also be interpreted as a special case of the Triple Momentum framework (7967721), which utilizes three momentum terms to achieve tighter global linear convergence bounds for deterministic, smooth, strongly convex functions compared to Nesterov’s method.
1:Initial iterate $x^{(1)}$, learning rate schedule $\gamma^{(t)}>0$, weight decay $\lambda\geq 0$, interpolation parameters $\mu_{x},\mu_{y}\in[0,1)$, base optimizer $\operatorname{\texttt{BaseOpt}}$.
[…]
3:for t=1,,Tt=1,...,T do
4:  $y^{(t)}=\mu_{y}x^{(t)}+(1-\mu_{y})z^{(t)}$ $\triangleright$ Update gradient computation point $y^{(t)}$.
5:  $g^{(t)}\in\partial f(y^{(t)};\xi^{(t)})$ $\triangleright$ Gradient is evaluated at $y^{(t)}$.
6:  $d^{(t)}=\operatorname{\texttt{BaseOpt}}(g^{(t)})$ $\triangleright$ Compute base optimizer’s search direction.
7:  $z^{(t+1)}=(1-\gamma^{(t)}\lambda)z^{(t)}+\gamma^{(t)}d^{(t)}$ $\triangleright$ Update $z^{(t)}$ iterate.
8:  $x^{(t+1)}=\mu_{x}x^{(t)}+\left(1-\mu_{x}\right)z^{(t+1)}$ $\triangleright$ Update weighted iterate average $x^{(t)}$.
9:end for
10:return $x^{(T)}$
Algorithm 1 Generalized Primal Averaging (GPA)
3.1 Interpreting GPA as Smoothed DiLoCo
As seen in Figure 1(a), increasing the number of inner steps leads to improved performance for single-worker DiLoCo. However, the underlying reasons for this behavior are not understood. By examining DiLoCo from the lens of GPA in Equation˜8 and comparing it with the more restrictive Nesterov formulation in Equation˜4, we can develop a deeper intuition for DiLoCo’s inner workings.
Suppose that we increase the number of inner steps in DiLoCo and want to maintain the same level of smoothing on the average iterate $x^{(t)}$. One may attempt to increase $\mu$ in Nesterov (Equation˜4) to decrease the weight on the current iterate $z^{(t+1)}$. However, since $\mu$ controls both the amount of smoothing in $x^{(t)}$ and the amount of interpolation used to update $y^{(t)}$, strictly increasing $\mu$ would decrease the recency of information from $z^{(t)}$ in $y^{(t)}$ by a factor of $\mu^{2}$, resulting in significantly different algorithmic behavior. Numerically, we validate that tuning $\mu$ alone in Nesterov’s primal averaging formulation is not sufficient to reach the performance of DiLoCo; see Appendix 11.
GPA addresses this limitation by decoupling the two roles of $\mu$ into separate hyperparameters: $\mu_{x}$ for the model evaluation sequence and $\mu_{y}$ for the gradient computation sequence. By controlling these two interpolation constants independently, we can smooth $x^{(t)}$ similarly without changing the amount of information introduced into $y^{(t)}$. This smoothing is depicted in Figure 2 on a simple deterministic quadratic problem. For a small number of inner steps, the methods closely align, but for a larger number of inner steps, their behavior diverges.
Figure 2: Comparison of DiLoCo and GPA’s trajectories on a deterministic quadratic problem. The outer iterates of DiLoCo are shown as red points, and the inner iterates as thin red lines.
Tuning GPA from DiLoCo. This intuition provides practical guidelines for converting a hyperparameter recipe for DiLoCo to GPA. Given an optimal number of inner steps $H$ and momentum parameter $\mu$ in DiLoCo, we observe for GPA that $x^{(t+H)}=\mu_{x}^{H}x^{(t)}+(1-\mu_{x})\sum_{k=0}^{H-1}\mu_{x}^{k}z^{(t+H-k)}$. Therefore, to match the coefficient in front of $x^{(t)}$ with DiLoCo, one can set $\mu_{x}=\mu^{1/H}$ while keeping $\mu_{y}\approx\mu$. With commonly used values $\mu=0.9$ and $H=32$, we obtain $\mu_{x}\approx 0.9967$ and $\mu_{y}\approx 0.9$. We leverage this heuristic to determine an effective number of inner steps used in Figure 1. See Table 6 for exact values mapping inner steps in DiLoCo to GPA coefficient $\mu_{x}$.
Tradeoffs with DiLoCo. GPA not only outperforms DiLoCo, but does so with fewer hyperparameters and lower memory requirements. While DiLoCo requires four hyperparameters, e.g., the inner and outer learning rate, momentum hyperparameter, and number of inner steps, the memory-efficient implementation of GPA reduces this to just three: the learning rate and two momentum parameters. This simplification is possible because DiLoCo’s practical performance is governed by an effective learning rate that couples the effect of the inner and outer learning rates ($\gamma^{(t)}$ and $\tilde{\gamma}$). On the other hand, GPA requires more FLOPs per-iteration, while DiLoCo amortizes its additional compute cost across multiple inner steps (see Appendix 9.3 for additional details).
4 Experiments
In this section, we assess the effectiveness of GPA on both language model pre-training and computer vision workloads. For language modeling, we compare against baselines AdamW, DiLoCo and Schedule-Free (SF), while for computer vision experiments, we compare GPA against AdamW. For DiLoCo, Schedule-Free and GPA methods, we use AdamW as the base optimizer (DiLoCo-AdamW, Schedule-Free and GPA-AdamW, respectively).
Table 2: Final validation loss versus effective number of inner steps $H$ for different optimizers on Llama-160M and Llama-1B models. Highlighted in bold is the lowest validation loss obtained across all inner step configurations $H$.
Llama-160M Llama-1B
Method $H=8$ $H=16$ $H=32$ $H=64$ $H=128$ $H=16$ $H=32$ $H=64$ $H=128$
[…]
(a) Comparison of AdamW, DiLoCo, and GPA with a fixed (effective) number of inner steps ($H=32$).
(b) Comparison of AdamW, DiLoCo, GPA and Schedule-Free with optimal hyperparameters.
Figure 3: Comparison of the validation loss against the number of steps for different optimizers on the Llama-160M workload.
4.1 Language Model Pre-Training
We conduct experiments on multiple scales of Llama models: 160 million and 1 billion parameters. These 160M and 1B models are pre-trained on the C4 dataset from scratch (2019t5) using a token budget of roughly 9.6 billion and 50 billion tokens, respectively (hoffmann2022training). All of our small experiments are conducted on two nodes equipped with two GB200 GPUs (184 GB of memory) while the larger scale model experiments utilize four nodes (with a total of 8 GPUs). Comprehensive details on batch size, sequence length, and hyperparameter sweeps can be found in Appendix 11. Note that both the Llama-160M and 1B experiments are performed in an overtrained setting.
Performance across number of inner steps. In Table 2, we provide the final validation loss for each method for different effective number of inner steps. Consistent with Figure 1(a), GPA-AdamW outperforms both DiLoCo-AdamW and AdamW across all settings of the effective number of inner steps.
Convergence behavior. Figure 3(a) shows the validation loss curves on Llama-160M for AdamW, DiLoCo-AdamW, and GPA-AdamW for the case where the number of inner steps is 32. In this case, $\mu_{x}$ has been tuned to match the number of inner steps; see Table 6 in Appendix 11 for details. In Figure 3(b), we compare GPA against the baselines (also including Schedule-Free-AdamW) by choosing the best performing runs over all hyperparameter choices including the effective number of inner steps. GPA-AdamW converges faster than both DiLoCo and AdamW throughout the entire training run. The training curves for GPA-AdamW are also noticeably smoother and more stable compared to the other methods. We observe that GPA-AdamW can handle higher learning rates compared to DiLoCo.
4.2 Code Generation Model Pre-Training
We also evaluate GPA on an 8 billion parameter Llama model designed for code generation; see Appendix 11.7 for more details. We find that GPA with any $\mu_{y}\in\{0.7,0.8,0.8\}$ outperforms AdamW throughout the course of training and achieves a better final validation loss (see Figure 4, which plots GPA with $\mu_{y}=0.7$).
Figure 4: Comparison of AdamW and GPA on the Llama-8B code generation model using 100 billion tokens.
4.3 Vision Transformer Model Training
To validate our method on a computer vision task, we train a ViT-S/16 model from timm on ImageNet with data augmentations from the repository. We train this under two batch size settings: (1) a small-batch setup with a batch size of 4,096 for 300 epochs; and (2) a large-batch setup with a batch size of 16,384 for 300 epochs. We tuned the methods separately in both settings, using the average over 2 random seeds to select the best hyperparameters, then ran the best-performing selection on 12 random seeds in total. For all methods, we used gradient clipping with norm 1 and a linear learning rate warmup over the first 5 epochs. We annealed the learning rate with cosine decay to $\times 0.001$ of the peak learning rate for all methods except Schedule-Free, which used constant learning rate after the warmup. Our results are that GPA outperforms AdamW by a clear margin in terms of validation accuracy throughout the course of training (see Figures 5). For further details on our hyperparameter tuning, see Appendix 11.
Figure 5: Comparison of AdamW, DiLoCo, and GPA on ImageNet ViT-S/16 from timm with data augmentations using a batch size of 4,096 samples.
5 Convergence Theory
Using the theoretical developments underpinning Schedule-Free learning, we can derive a convergence bound for GPA given any base optimizer that has a regret bound, using the framework of online-to-batch conversion (cesa2004generalization). We will use the Bregman divergence of $F$ defined as $B_{F}(a,b)=F(a)-F(b)-\langle\nabla F(b),a-b\rangle$ for $a,b\in\mathbb{R}^{n}$.
Theorem 1.
Let $F$ be a convex function and assume that there exists a minimizer $x_{*}$ that minimizes $F$. Let $\xi^{(1)},\dots,\xi^{(T)}$ be a sequence of i.i.d. random variables. Suppose that we are given arbitrary updates $z^{(1)},\dots,z^{(T)}$ from a base optimizer within the Generalized Primal Averaging framework (Equation 8). Then for $\mu_{x},\mu_{y}\in[0,1)$ and average iterate $\bar{x}^{(T)}=\frac{1}{T}\sum_{t=1}^{T}x^{(t)}$, we have the bound
$\mathbb{E}[F(\bar{x}^{(T)})-F(x_{*})]$ $\leq\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\langle\nabla F(y^{(t)}),z^{(t)}-x_{*}\rangle]+\frac{\mu_{x}}{1-\mu_{x}}\frac{1}{T}\mathbb{E}\left[F(x^{(1)})-F(x_{*})\right]$
$\qquad-\frac{1}{1-\mu_{y}}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[B_{F}(y^{(t)},x^{(t)})]-\frac{\mu_{y}}{1-\mu_{y}}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[B_{F}(x^{(t)},y^{(t)})]$
$\qquad-\frac{\mu_{x}}{1-\mu_{x}}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[B_{F}(x^{(t-1)},x^{(t)})].$
Corollary 1.
Assume that the base optimizer has the regret guarantee $\sum_{t=1}^{T}\mathbb{E}[\langle\nabla F(y^{(t)}),z^{(t)}-x_{*}\rangle]=\mathcal{O}(\sqrt{T})$. Then:
$\mathbb{E}[F(\bar{x}^{(T)})-F(x_{*})]=\mathcal{O}\left(\frac{1}{\sqrt{T}}\right).$
Remarks on Theorem 1:
  • The first term on the right-hand side of the regret bound is the average regret of the base optimizer. This term captures the convergence rate from the base optimizer.
  • The second term has a positive term, which decays at a rate of $1/T$, which is typically faster than the decay of the term in the first row.
  • All remaining Bregman divergence terms are negative, and so are potentially beneficial. If $\mu_{x}$ and $\mu_{y}$ are chosen such that the negative terms dominate the positive second term, then GPA will converge faster than the base optimizer.
  • The same terms appear in the convergence guarantees for Schedule-Free methods, and can explain when they may work better. For strongly convex problems, such Bregman divergences were used to get $\mathcal{O}(1/T)$ convergence.
  • Unlike the guarantees for Schedule-Free, our convergence bound is for the average iterate. For best performance, a learning rate schedule should be used and the last iterate returned (defazio2023when).
  • Our bound indicates that GPA will be faster than the base optimizer when the objective function varies nonlinearly between consecutive iterates and between $x^{(t)}$ and $y^{(t)}$.
6 Conclusion
GPA generalizes Nesterov momentum by decoupling interpolation constants for gradient computation and model evaluation. On both small and large-scale language models, GPA outperforms single-worker DiLoCo while eliminating its complex two-loop structure, simplifying hyperparameter tuning, and reducing memory overhead.
[…]
Finally, GPA’s decoupling of parameters also enables new avenues for distributed training. In DiLoCo, the number of inner steps serves as a coupled hyperparameter for both Step-$K$ Nesterov and local SGD, leading to the undesirable finding that increasing the number of inner steps can improve convergence – contrary to standard local SGD intuition. By introducing a tunable, continuous smoothing parameter that is independent of the number of inner steps in Step-$K$ Nesterov, GPA provides a new foundation for re-designing algorithms for cross-regional training.
Acknowledgements
We thank Anna Cai, Jiaming Cui, Runa Eschenhagen, Shagun Gupta, Yuchen Hao, Chien-Chin Huang, Minhui Huang, Dzmitry Huba, Vladimir Ivanov, Dominik Kallusky, Tsung-Hsien Lee, Vishal Nandavanam, Shangfu Peng, Vinay Rao, Junjie Wang, Gavin Zhang, Iris Zhang, Xin Zhang, and Chuanhao Zhuge for their discussions on the algorithm, support on experimentation in TorchTitan, and detailed feedback on the manuscript. We also thank Kristin Lauter, Maxim Naumov, Sandeep Parab, Joel Pobar, and Chunqiang Tang for their managerial support of this work. Finally, we thank Derek Shao, Emrah Seker and Rodrigo Paim from the Meta Superintelligence Labs Infrastructure team for supporting us with torchtitan conda packages in a timely manner to enable faster experimentation.